OpenStudy (anonymous):

m<1=75, m<2=110, m<3=100

OpenStudy (vishweshshrimali5):

do u want to calculate value of m ?

jhonyy9 (jhonyy9):

m-1<=75 m-2<=110 m-3<=100 - you have thought the same ?